I just recently have learned how to find the vertex of a function. However, with this problem, the fraction ... confused me. :) I get half way through, then get entirely lost!


Determine the vertex of each function by completing the square.

y= 8x^2 + -14x + -72

Thanks!
-B

Answer by stanbon(48558)

Determine the vertex of each function by completing the square.
y= 8x^2 + -14x + -72
-------------------------
Rewrite:
8x^2 - 14x = y + 72
---
8(x^2 - (7/4)x + ?) = y + 8?
Complete the square:
--
8(x^2 - (7/4)x + (7/8)^2)) = y + 8(7/8)^2
---------------------
8(x-(7/8))^2 = y + +8*(49/64)
---
8(x-(7/8))^2 = y + (49/8)
-------------------------------
Vertex: ((7/8) , (-49/8))
==============================
Cheers,
Stan H.